OpenStudy (anonymous):

f(x)=x^2*e^(-|x|) how can i solve the integral?

OpenStudy (amistre64):

i now there are some representations of e^x that cant be integrated.... forget why..

OpenStudy (amistre64):

but for this; just int by parts using x^2 as u; and v as e^...

OpenStudy (anonymous):

Yes, but how i deal the absolute value?

OpenStudy (amistre64):

an absolute value is always positive; and the opposite of an always positive value is always negative right?

OpenStudy (amistre64):

just redefine it; |x| = a perhaps

OpenStudy (anonymous):

Shouldn't i divide the integral into two cases? |x| = x if x>0 and |x| = -x if x<0?

OpenStudy (shadowfiend):

Yeah, I think that's the way you have to go.

OpenStudy (watchmath):

Is this a definite integral or indefinite?

OpenStudy (anonymous):

The exercise is to solve the integral from -inf to inf

OpenStudy (watchmath):

That makes a difference \(\int_{-\infty}^\infty x^2e^{-|x|}=2\int_0^\infty x^2e^{-x}\,dx\) and you can use integration by parts.

OpenStudy (amistre64):

thats what I had in mind :) just couldnt spell it out well enough ...